Anonymous1babes:

Solve: negative 1 over 2 x + 3 = −x + 7 (5 points) a 1 over 2 b 8 c −8 d 2

juju2021:

ummm

Sailor:

Ok.. So you would first add x to both sides to cancel out -x.

juju2021:

subtract 3 from both sides

Sailor:

Then divide by 3

Anonymous1babes:

okay so that would give me ?

Sailor:

Do 4 divided by 3.

Sailor:

Wait- that's not an answer choice

juju2021:

its not

Sailor:

@juju2021 wrote:
its not
Maybe we did something wrong?

snowflake0531:

\[-\frac{1}{2}x+3=-x+7 \\ -\frac{1}{2}x+3+x=-x+x+7\\\frac{1}{2}x+3=7\\\frac{1}{2}x+3-3=7-3\\\frac{1}{2}x=4\\\frac{1}{2}x(2)=4(2)\\x=8\]

Sailor:

@snowflake0531 wrote:
\[-\frac{1}{2}x+3=-x+7 \\ -\frac{1}{2}x+3+x=-x+x+7\\\frac{1}{2}x+3=7\\\frac{1}{2}x+3-3=7-3\\\frac{1}{2}x=4\\\frac{1}{2}x(2)=4(2)\\x=8\]
:0 ah i see now

juju2021:

u cant do that

Sailor:

Nono, we got something wrong

juju2021:

ik we did but what

Sailor:

It was 1/2x not 2x

snowflake0531:

ahem it was -1/2 x

juju2021:

i see it now

Sailor:

🤦 Thanks for helping snowflake!

juju2021:

yea thanks

Anonymous1babes:

thank you guys for helping me

juju2021:

np

Olivia144:

@juju2021heeiyy